• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Anklin

    Log entry time 10:46:06 on February 24,1998

    Entry number 58

    Followups:

    keyword=beam current analysis

    BCM data for the relevant runs are shown in the next few plots. The analys is based on the 16707 calibration.

    Figure 1:The Ratio of BCM1 / BCM2 starts to be beam current dependent at about
    run 17050. The same behaviuor can also be seen in 1/2 and 2/3. That means
    more than one BCM drifted ! As a first assumption this could be 1 and 2,
    since they are located in the same crate. That crate's power supply turned
    bad and has been replaced later during the thanksgiving break. With the
    intrinsic data (charge#####.txt) further studies can be done. Either the
    offsets or the gains could have drifted!!!
    At run 17131 two things have changed: first the beam current dropped
    from roughly 100 microA to 40 microA and the unused frequency generator
    has been removed from the faky crate.
    After the thanksgiving break things look more constant. Beam current
    was kept constant. The dip in the 1/2 ratio has already been explained in
    hclog entry 2788. it is very likely due to a instability of BCM1.

    Figure2: The expected (!) current dependence of the ratio 1/2 from calibration
    run 16707. Due to different gain settings of 1 and 2 a nonlinearity
    is expected at very low beam currents, where the one with the higher
    gain starts to fall of the linear range of the amplifier.
    But the expected deviation at 30 microA is about 0.05% !!!

    The plots 1/3, 2/3 and beam current monitored with bcm2 are shown in the following
    fpilog entries:


    FIGURE 1

    FIGURE 2